OpenStudy (pokiedokie):

I have to rewrite these equations in slope-intercept form, am I correct? -4x + 6y = 18 and 3y + x = 5

OpenStudy (pokiedokie):

for -4x + 6y = 18 I rewrote it to: y = 18 - 4x + 6y

OpenStudy (pokiedokie):

I'm about 99.9% sure I got that incorrect, can someone help?

OpenStudy (ckallerid):

It's a bit off. If I remember correctly in slope intercept form you're supposed to isolate the y. Do you know what PEMDAS is?

OpenStudy (pokiedokie):

yes i do, so I solve y first and use PEMDAS?

OpenStudy (ckallerid):

well kinda... You need to be doing both at the same time.

OpenStudy (pokiedokie):

that's what i meant aaah sorry

OpenStudy (ckallerid):

Its alright lol

OpenStudy (ckallerid):

so first just start by adding 4x to both sides It will come out as \[6y = 4x +18\]

OpenStudy (ckallerid):

Divide 6 to both sides \[y = \frac{ 4x }{ 6 } + \frac{ 18 }{ 6 }\]

OpenStudy (ckallerid):

simplify \[y = \frac{ 4x }{ 6 } + 3\] And you have that answer

OpenStudy (pokiedokie):

oh okay! so we leave the fraction as is then?

OpenStudy (ckallerid):

yup

OpenStudy (ckallerid):

same with the other one. 3y + x = 5 Subtract x from both sides 3y = -x + 5 Divide 3 to both sides y = -x/3 + 5/3 You can't simplify on this one sooooo that's the answer!

OpenStudy (pokiedokie):

thank you so much!! i tried the 2nd one and ended up getting it correct like you said. I understand it much better now ty!